OpenStudy (anonymous):

Find the work in ft-lbs required to empty a full hemispherical tank with a radius of 6 ft by pumping the water to 2 ft above the top of the tank

OpenStudy (anonymous):

@KingGeorge

OpenStudy (kinggeorge):

I'll be totally honest. I forgotten all the details of how to do this kind of problem. I do know that\[W=\int F\,dx\]where W is work, F is force, and x is some distance. I also remember that in your case, the integral will be from 2 (2 feet above the tank) to 8 (radius + 2).

OpenStudy (anonymous):

I have some notes

OpenStudy (anonymous):

this is for a similar problem

OpenStudy (kinggeorge):

What's in front of the integrals? A "D"?

OpenStudy (anonymous):

Yeah, those are Ds

OpenStudy (kinggeorge):

What do they stand for?

OpenStudy (anonymous):

hero made the notes

OpenStudy (anonymous):

my guess is distance

OpenStudy (kinggeorge):

Alright, I think I've got a little better idea of what to do know. D stands for density, and g is the gravitational constant. The density is required because we also need to factor in the mass, and mass=density*volume. So the only thing I'm not sure about at this point is the volume.

OpenStudy (anonymous):

One minute.

OpenStudy (anonymous):

v = The Integral of f(x) dx from b to a. I think... I'm on my phone right now. The formula was drawn out in another post

OpenStudy (kinggeorge):

I feel a little bad about posting a link, but I think there is a very clear explanation of a very similar problem here. Simply change the units/constants, and the only thing left, is that you have to lift everything an extra 2 feet. http://www.math.washington.edu/~yumiko88/Final/Work/F%20Work%20Sol.pdf

OpenStudy (anonymous):

Will you help for this one so I have an example to go off of?

OpenStudy (kinggeorge):

Sure. I'll be basing my explanation off of what's in the link though.

OpenStudy (anonymous):

Thank you. I can't even click the link on mobile. I'll get on my computer soon though

OpenStudy (kinggeorge):

First, I'm going to model it with a graph of a semi-circle that's centered at (0,-2) (so we have to lift an extra 2 units). The equation for the circle is\[x^2+(y+2)^2=36\]and we're taking the lower half, so the equation reduces to\[x=\sqrt{32-4y-y^2}\]Now we find the volume.

OpenStudy (kinggeorge):

Since \(x\) is modeling the radius, the volume at some point \(y\) is simply\[\pi x^2\,\Delta y=\pi(32-4y-y^2)\Delta y\;ft.^3.\]Where the \(\Delta y\) is the height. From this, we can find the mass. The mass is simply density*volume, and for water, we know the density to be \(62.4 \;lbs./ft.^3\). So the mass is\[62.4\pi(32-4y-y^2)\Delta y\;lbs.\] Next is the force term.

OpenStudy (kinggeorge):

But this is also easy, since the force is merely mass*acceleration, and in our case, I think the constant is usually just \(32 \;ft/s^2.\) So the force is\[32\cdot62.4\pi(32-4y-y^2)\Delta y\;lbs\cdot ft/s^2\]

OpenStudy (kinggeorge):

Lastly, we can't forget that we have to move the water up \(-y\) units (since we're assuming everything is below the x-axis). So we have to multiply by \(-y\) as well. Finally, we integrate from -8 to -2. So if everything works out, it should be that the work is\[-\int_{-8}^{-2}32\cdot62.4\cdot\pi\cdot y\,(32-4y-y^2)\,dy\]

OpenStudy (anonymous):

I will test that on my computer and give you what I got

OpenStudy (kinggeorge):

I really hope that's right...

OpenStudy (kinggeorge):

So I just discovered a slight error (it doesn't effect the actual integral though). When I first said the equation reduced to \[x=\sqrt{32-4y-y^2}\]it should have been\[x=-\sqrt{32-4y-y^2}\]instead. But since we immediately squared this to get the volume, the negative sign doesn't matter.

OpenStudy (anonymous):

This is odd. From my knowledge, the final answer is in Pi. Ex.: 9pi

OpenStudy (kinggeorge):

There is a \(\pi\) in my integral. Which you can factor out if necessary since it's a constant.

OpenStudy (kinggeorge):

So if you want, you could write it as\[-\,\pi\int_{-8}^{-2}32\cdot62.4\cdot y\,(32-4y-y^2)\,dy\]

OpenStudy (anonymous):

I'm on my computer. I'll try that.